Required information Exercise 3-14A Prepare an adjusted trial balance (LO3-3, 3-4) The following information applies to the questions displayed below. The December 31, 2021, unadjusted trial balance for Demon Deacons Corporation is presented below. Debit $ 8,400 13,400 5,280 2,400 Accounts Credit Cash Accounts Receivable Prepaid Rent Supplies Deferred Revenue Common Stock Retained Earnings Service Revenue Salaries Expense 1,400 11,000 4,400 39, 680 27,000 $56,480 $56,480 At year-end, the following additional information is available 1. The balance of Prepaid Rent, $5,280, represents payment on October 31, 2021, for rent from November 1, 2021, 2. The balance of Deferred Revenue, $1,400, represents payment in advance from a customer. By the end of the . An additional $700 in salaries is owed to employees at the end of the year but will not be paid until January 4 4. The balance of Supplies, $2,400, represents the amount of office supplies on hand at the beginning of the year to April 30, 2022 year, $350 of the services have been provided. 2022 of $900 plus an additional $1,500 purchased throughout 2021. By the end of 2021, only $640 of supplies remains.
Exercise 3-14A Part 2 2. Prepare an adjusted trial balance as of December 31, 2021 DEMON DEACONS CORPORATION Adjusted Trial Balance December 31, 2021 Accounts Debit Credit Cash Accounts Receivable Prepaid Rent Supplies Deferred Revenue Salaries Payable Common Stock Retained Earnings Service Revenue Salaries Expense Rent Expense Supplies Expense Totals 0
Answer #1
Adjusted Trial balance
December 31,2021
Debit Credit
Cash 8400
Accounts Receivable 13400
Prepaid Rent 3340 =5280-1940
Supplies 640
Deferred Revenue 1050 =1400-350
Salaries Payable 700
Common Stock 11000
Retained Earnings 4400
Service Revenue 40030 =39680+350
Salaries expense 27700 =27000+700
Rent expense 1940 =5820*2/6
Supplies expense 1760 =2400-640
Totals 57180 57180